OpenStudy (anonymous):

find the points of intersection of the parabolas y=1/2x^2 and y=1-1/2x^2. Show that at each of these points the tangent lines to the 2 parabolas are perpendicular. Thanks!

OpenStudy (anonymous):

the intersection when y =1/2x^2=y=1-1/2x^2

OpenStudy (anonymous):

tangent means the slope which is the derivative of the function

OpenStudy (anonymous):

pertpendicular..... slope1*slope2=-1

OpenStudy (anonymous):

i dont know what is a derivative. i just start calculus this semester, and we're starting on change of limits.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

plug x value (intersection) to the function y=1/2 x^2

OpenStudy (anonymous):

i dont sure where to get the x value

OpenStudy (anonymous):

m = \[m = \frac{ Y _{2} - Y _{1} }{ X _{2} - X _{1} }\]

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

The best thing you can do is to go Read that section in your book, carefully. That's what I always do, and find the answers for myself.

OpenStudy (campbell_st):

I think the 1st task should be to find the points of intersection so \[\frac{1}{2} x^2 = 1 - \frac{1}{2} x^2\] just so you know the points you need.

OpenStudy (campbell_st):

and I'd recommend graphing the 2 curves... to help you understand what is happening...

OpenStudy (anonymous):

okay...

OpenStudy (sissyedgar):

... :)